# Differential Test Performance and Peer Effects

### v1

@@ 2.3 Identification
where $z$ denotes variables excluded from $X$ that may be used as instruments. Let $\mathcal{Z}_{n}^{*}=\sigma\left(S,\alpha,X,z,\zeta^{\ast}\right)$ be the sigma field of all conditioning variables and let $\mathcal{Z}_{n}=\sigma\left(X,z\right)\subset\mathcal{Z}_{n}^{*}$ be the subset of observable information.
The information set $\mathcal{Z}_{n}^{*}$ consists both of observable characteristics $X,$ classroom allocations $S,$ excluded variables $z$ that may be used as instruments as well as of unobserved variables $\zeta^{\ast}$ and $\alpha.$ The following proportionality restriction on the conditional means of the latent outcomes is at the core of our proposed method of identifying peer effects.
Assumption 2.
Assume that
$E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]=\kappa_{i}f_{t}$
where $\kappa_{i}$ is a random variable that is invariant to $t$ and is measurable with respect to $\mathcal{Z}_{n}^{*}$ and $f_{t}$ is a fixed parameter that only varies with $t$. In addition, assume that the random variable $y_{it}^{*}$ is either bounded, or there exists an $\eta>0$ and a random variable $y$ such that $\left|y_{it}^{*}\right|+\left|\kappa_{i}f_{t}\right|\leq y$ with $E\left[\left|y\right|^{4+\eta}|\mathcal{Z}_{n}^{*}\right]\leq K_{y}<\infty$ for all $i$ and $t.$ In addition, $E\left[\left|\alpha_{c}\right|^{2+\eta}\right]\leq K_{\alpha}<\infty$ for all $c.$
Without loss of generality we normalize $f_{2}=1$. The assumption includes the special case where $f_{t}=1$ for $t=1,2$, which is relevant in situations where tests $t$ measure closely related skills. The interpretation of the condition is that conditional on a hypothetical or actual baseline of test results $\zeta^{*}$, as well as information about group formation $S$, unobserved class level characteristics $\alpha$, as well as other observed characteristics $X$ and $z$, additional test results do not change expected performance, except maybe for a common scale factor $f_{t}$ accounting for systematic differences between tests affecting all test takers in the same fashion. [^4] For simplicity we assume that $f_{t}$ in Assumptions 1 and 2 are the same. This restriction can be relaxed if three test outcomes are used. The factors $f_{t}$ can be treated as random rather than fixed at the expense of slightly more complex asymptotic arguments needed for the analysis of statistical inference, see Kuersteiner and Prucha (2020). It is in this sense that tests $t=1$ and $t=2$ measure essentially identical skills. Note that $E\left[y_{it}^{*}|\mathcal{Z}_{n}^{*}\right]=\kappa_{i}f_{t}$ by iterated expectations and $\mathcal{Z}_{n}^{*}\subset\mathcal{F}_{n,i,t}.$
The variable $\kappa_{i}$ is a function of the entire cross section of latent performance measures $\zeta^{*}$, of the allocation $S$ and observed and unobserved characteristics. It is generally cross-sectionally dependent in ways we do not restrict or specify and captures such common effects as the socioeconomic background of students and classmates, their family background including parental education and support, their interaction with peers, the exposure to their teachers and resources available in their classroom. It also depends on a student’s own ability. The restriction we impose is that these factors do not change between tests $t=1$ and $t=2$, except maybe for a change in scale $f_{t}.$
Baseline performance $\zeta^{*}$ does not necessarily have to be observed or realized and $\zeta^{*}$ can alternatively be interpreted as latent ability prior to taking the actual tests $t=1,2$. In this sense, $\kappa_{i}f_{t}$ and $E\left[y_{it}^{*}|\mathcal{Z}_{n}^{*}\right]$ are also unobserved. A scenario where Assumption 2 is realistic arises when the same type of test is taken multiple times without additional training between iterations of the test, or when several tests are given that focus on related skills, as is the case in Project STAR data for grades K-3 and tests for reading, word study and listening skills. [^5] Another example of a test with multiple scores is the Comprehensive Testing Program (CTP) administered by the Educational Records Bureau and given to Grades 1-11. The CTP test consists of several main categories such as “Auditory Comprehension”, “Reading Comprehension” and “Mathematics” with each category consisting of additional subcategories that each receive separate scores. The assumption is less plausible if comparisons are attempted across different subjects in a high school or college setting, for tests given at different times during the school year, or measuring different skills in a professional or team setting. In the latter case, additional factors can be introduced if multiple measurements per skill category are available. [^6] The multi-factor case is discussed in detail in Kuersteiner and Prucha (2020).
The following theorem formally establishes a representation for $y_{it}^{*}$ that decomposes $y_{it}^{*}$ into its conditional mean, $E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]$, that is dependent cross-sectionally as well as across tests, and into uncorrelated idiosyncratic noise $u_{it}$. The theorem formalizes the definition of $u_{it}$
Theorem 2.1.
Suppose Assumption 2 holds. Define
$u_{it}=y_{it}^{*}-E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]=y_{it}^{*}-\kappa_{i}f_{t},$
then by construction
$E\left[u_{it}|\mathcal{F}_{n,i,t}\right]=0\ \ \text{and}\ \ \operatorname{Cov}\left(u_{it}u_{js}|\mathcal{F}_{n,i,t}\right)=0$
for any $j\neq i$ or $t\neq s$. Furthermore, let $\mathcal{G}_{n,i,t}$ be any sigma field with $\mathcal{G}_{n,i,t}\subseteq\mathcal{F}_{n,i,t}$, then $E\left[u_{it}|\mathcal{G}_{n,i,t}\right]=0$ and $\operatorname{Cov}\left(u_{it},u_{js}|\mathcal{G}_{n,i,t}\right)=0$ for any $j\neq i$ or $t\neq s$, and $\operatorname{Cov}\left(u_{it},\kappa_{i}f_{t}\right)=0,\operatorname{Cov}\left(u_{it},X\right)=0$, $\operatorname{Cov}\left(u_{it},z\right)=0$ etc. If $E\left[y_{it}^{*}-\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]\neq 0$ for some $\mathcal{G}_{n,i,t}$, then Assumption 2 cannot hold.
Examples for the information sets $\mathcal{G}_{n,i,t}$ in the above theorem are $\mathcal{G}_{n,i,t}=\sigma\left(S,\alpha,X,z,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*s}\right)$ where $y_{-i,t}^{*s}$ is a subvector of $y_{-i,t}^{*}$ or $\mathcal{G}_{n,i,t}=\mathcal{Z}_{n}^{*}.$ The above theorem makes clear that the main restriction we impose on $y_{it}^{*}$ is the assumption that latent test performance is not predictable conditional on observable characteristics and unobserved baseline ability. The properties $E\left[u_{it}|\mathcal{Z}_{n}^{*}\right]=0,$ $\operatorname{Cov}\left(u_{it},u_{js}|\mathcal{Z}_{n}^{*}\right)=0$ for any $j\neq i$ or $t\neq s$ and $\operatorname{Cov}\left(u_{it},\kappa_{i}f_{t}\right)=0$, $\operatorname{Cov}\left(u_{it},X\right)=0$, $\operatorname{Cov}\left(u_{it},z\right)=0$ follow directly from Assumption 2 and are not additional assumptions imposed on the distribution of $y_{it}^{*}.$ In other words, Assumption 2 is necessary and sufficient for the decomposition of $y_{it}^{*}=\kappa_{i}f_{t}+u_{it}$ with uncorrelated errors $u_{it}$.
We do not attach a specific economic or behavioral interpretation to the decomposition of $y_{it}^{*}$ into $\kappa_{i}f_{t}$ and $u_{it}$. It is merely a statistical representation for the purpose of isolating variation in test scores that is correlated within groups but invariant across tests and variation that is idiosyncratic to the individual student and test. Normalizing $f_{2}=1$, which is without loss of generality, leads to an immediate consequence of the decomposition: The quasi-difference
$y_{i1}^{*}-y_{i2}^{*}f_{1}=u_{i1}-u_{i2}f_{1}$ (3)
is mean zero and uncorrelated across $i$ conditional on $\mathcal{Z}_{n}^{*}$. We exploit these two moment restrictions to formulate our GMM estimator. They are at the heart of our identification result which is formally presented in Section 4.
Our framework is in contrast with much of the econometrics literature on the identification of peer effects where assumptions about the orthogonality of $u_{it}$ and $\kappa_{i}f_{t}$ and covariates are usually imposed directly. These unobservable quantities often are given economic interpretations. Examples of papers with such assumptions include Lee (2007b) who assumes i.i.d. errors, Graham (2008) who assumes independent classroom and idiosyncratic errors conditional on class type, Arcidiacono et al. (2012) who assume uncorrelated idiosyncratic errors and fixed effects that are orthogonal to idiosyncratic errors, Rose (2017) who assumes uncorrelated idiosyncratic errors, Kuersteiner, Prucha, and Zeng (2023) who assume a random effects specification and Lewbel, Qu, and Tang (2023) who assume exogeneity of group formation. Often these assumptions are justified by random group selection. In our setting we neither assume that individuals are randomly assigned to groups, nor that they are randomly selected from a population or that they have otherwise observed or unobserved characteristics that are independent in the cross-section.
2.4 Estimation
Our estimators are using differential, or quasi-differential test scores, $y_{i1}-y_{i2}f_{1}$. These differential measures purge performance measures from common unobserved influences captured by $\alpha_{c}$ and $\kappa_{c}=\left(\kappa_{c1},\ldots.\kappa_{cn_{c}}\right)^{\prime}=S_{c}\tilde{\kappa}$ and shed light on the existence of peer effects. Define the unobserved class and individual effect as
$\mu_{c}^{\ast}=\alpha_{c}\mathbf{1}_{c}+\left(I_{c}+\rho M_{c}\right)\kappa_{c}$ (4)
such that Assumptions 1, 2 and Theorem 2.1 imply the empirical specification
$y_{ct}$ $=\mu_{c}^{*}f_{t}$ $+v_{c}^{c}\beta_{vt}^{c}+w_{ct}^{c}\beta_{wt}^{c}+\left(I_{c}+\rho M_{c}\right)(v_{c}^{p}\beta_{vt}^{p}+w_{ct}^{p}\beta_{wt}^{p}+u_{ct}).$ (5)
By quasi-differencing (5), which eliminates classroom and individual fixed effects, we obtain
@@ 2.4 Estimation
We treat $f_{t}$ as an unknown parameter to be estimated that accounts for differences in average test scores between tests $1$ and $2$. Collecting terms with $f_{1}$ leads to an interpretation of (2.4) where $y_{c2}-v_{c}^{c}\beta_{v2}^{c}-w_{c2}^{c}\beta_{w2}^{c}-\left(I_{c}+\rho M_{c}\right)(v_{c}^{p}\beta_{v2}^{p}+w_{c2}^{p}\beta_{w2}^{p})$ is used to control for unobserved $\mu_{c}^{\ast}$ similar to a control function approach. The difference to a conventional control function approach is that (5) for $t=2$ cannot be consistently estimated by least squares because unobserved components may be correlated with observed covariates, a problem that our GMM estimators address.
We next write (2.4) more compactly by stacking observations across classrooms. Let $y_{t}$ , $v^{p}$ and $w_{t}^{p}$ be as defined earlier. In similar fashion define $v^{c}$, $w_{t}^{c}$ and $u_{t}$, and let $M=\operatorname{diag}_{c=1}^{C}\left(M_{c}\right)$. Let $\underline{X}=\left(v^{c},w_{1}^{c},w_{2}^{c},v^{p},w_{1}^{p},w_{2}^{p},Mv^{p},Mw_{1}^{p},Mw_{2}^{p}\right)$ and define
@@
where $\delta_{v}^{c}=\beta_{v1}^{c}-f_{1}\beta_{v2}^{c}$, $\delta_{v}^{p}=\beta_{v1}^{p}-f_{1}\beta_{v2}^{p}$ and $\beta$ is the vector of all $\beta$s above, then we can write (2.4) for the entire sample more compactly as
$y_{1}=f_{1}y_{2}+\underline{X}\delta+(I+\rho M)(u_{1}-f_{1}u_{2})$ (8)
@@
where $\delta=\beta_{1}-f_{1}\beta_{2}$ with $\beta_{t}$ collecting the corresponding coefficients for each variable in $\underline{X}$ and set to 0 if a variable is absent in (5) for given $t$. Similarly we can define $\mu^{\ast}=(\mu_{1}^{\ast\prime},....,\mu_{C}^{\ast\prime})^{\prime}$ and write (5) compactly as
$y_{t}=\mu^{\ast}f_{t}+\text{$\underline{X}$$\beta_{t}+(I+\rho M)u_{t}$},$ (9)
@@
With Assumptions 1 and 2, and standard regularity conditions on $\underline{X}$, we can identify $\delta$, which is sufficient for the identification of $(\rho,f_{1})$. The parameters $\beta_{w1}^{c}$ and $\beta_{w1}^{p}$ can be directly recovered from $\delta$, and $\beta_{w2}^{c}$ and $\beta_{w2}^{p}$ can be identified when $f_{1}\neq 0$. As is common in pure fixed effect panel settings, the effects of test invariant covariates $v$, in our case $\beta_{v1}^{c},\beta_{v2}^{c}$ and $\beta_{v1}^{p},\beta_{v2}^{p}$ are not identified when using within or differencing estimators. In our case, except when $f_{1}=1,$ we are quasi-differencing the equation and allowing for heterogeneity in parameters between $t=1$ and $t=2.$ This leads to a formulation of the model where $v^{c}$ and $v^{p}$ enter the quasi, or fully differenced equation. However, without further assumptions the parameters $\beta_{v1}^{c}$, $\beta_{v2}^{c}$, $\beta_{v1}^{p}$ and $\beta_{v2}^{p}$ are not separately identified. An example of restrictions where the parameters $\beta_{v1}^{c}$ and $\beta_{v2}^{c}$ are identified arises when $f_{1}\neq 1$ and $\beta_{v1}^{c}=\beta_{v2}^{c}$.
@@
In empirical applications it is sometimes sufficient to estimate $\beta$ or even just $\delta,$ for example when $\underline{X}$ contains contextual peer effects. Conventional regression methods applied directly to (5) are invalid because unobserved fixed effects collected in $\mu_{c}^{\ast}$ may be correlated with $\underline{X}.$ We show in Section 4 that instrumenting for $y_{2}$ in (8) identifies the parameters $\delta$ and $f_{1}.$ We show in Corollary 4.1 that when test scores are recorded as non-negative values the classroom fixed effect or an overall constant is a valid instrument. More generally, exogenous covariates that vary at the classroom level but are excluded from $\underline{X}$ are valid instruments by Lemma 4.1.
Valid standard errors are obtained by clustering by classrooms. For estimation of the full set of parameters including the peer effect parameter $\rho$, we use the efficient GMM estimator developed in Kuersteiner and Prucha (2020) for a fairly general class of networks covering both social and spatial network effects. Identification of the parameters $\rho,f_{1}$ and $\delta$ follows from Assumptions 1 and 2 as well as classroom level homoskedasticity and some restrictions on the parameter space that are detailed in Section 4.
Feasible efficient inference requires some additional restrictions on conditional variances that we now describe. We allow for some forms of heteroskedasticity for $u_{it}$ across tests and by class type. Let $\tau_{c}$ be a categorical variable for class type with $\tau_{c}=j$ if classroom $c$ is of type $j\in\left\{1,...,J\right\}.$ Examples of types are small and regular classes, or classes located in urban, suburban or rural areas.
Assumption 3.
Let $u_{it}$ be as defined in Theorem 2.1 and let $i=n_{1}+\ldots+n_{c-1}+r$ with $1\leq r\leq n_{c}$, i.e., $i$ refers to student $r$ in class c. Let $u_{-i,t}=\left(u_{1,t},...,u_{i-1,t},u_{i+1,t},...,u_{n,t}\right)$. Define the filtration
$\mathcal{B}_{n,i,t}=\sigma\left(S,\alpha,X,z,\zeta^{*},u_{-t},u_{-i,t}\right)\text{\,for $t=1,2$}$
where $u_{-t}$ represents information about the other test. Then the conditional variance of $u_{it}$ is given by
$E\left[u_{it}^{2}|\mathcal{B}_{n,i,t}\right]=\sigma_{t}^{2}\rho_{\tau_{c}}^{2},$
where $\tau_{c}$ is $\mathcal{Z}_{n}$-measurable, and $c_{u}\leq\sigma_{t}^{2},\rho_{j}^{2}\leq C_{u}$ for some finite positive bounds.
Assumption 3 implies that for $t=1,2$, $\Sigma_{t}\equiv E\left[u_{t}u_{t}^{\prime}|\mathcal{Z}_{n}^{\ast}\right]=\sigma_{t}^{2}\operatorname{diag}_{c=1}^{C}\left(\rho_{\tau_{c}}^{2}I_{c}\right)$. Observing that $f_{t}$ is a fixed constant, conditional on $S,\zeta^{*},\alpha,X$ and $z$, knowledge of $y_{it}^{*}$ is equivalent with knowledge of $u_{it}$. Thus we have the following equivalence of information sets [^7] See the proof of Lemma B.3 for more detail.
=$\mathcal{F}{}_{n,i,t}=\mathcal{B}_{n,i,t}\,\text{for}\,t=1,2.$
Assumption 3 imposes a similar factor structure on conditional variances as does Assumption 2 for conditional means.
Below we illustrate the nature of our efficient quasi-differencing transformation. By Theorem 2.1 the errors $u_{it}$ are mutually uncorrelated over tests as well as cross-sectionally. Consequently the variance covariance matrix of $u_{1}-f_{1}u_{2}$ is given by,
Define $\gamma_{j}^{2}=\left(\sigma_{1}^{2}+f_{1}^{2}\sigma_{2}^{2}\right)\rho_{j}^{2}$, for $j=1,...,J$ and $\gamma=\left(\gamma_{1},...,\gamma_{J}\right)^{\prime}$. Then $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}\left(\gamma_{\tau_{c}}^{2}I_{c}\right).$
The disturbance term in (8) is of the form $(I+\rho M)(u_{1}-f_{1}u_{2})$. Denote the parameter $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$. Premultiplying (8) by $\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}$ then yields the following model
$y^{+}(\theta,\gamma)=X^{+}(\theta,\gamma)\delta+u^{+}$ (10)
where,
$y^{+}(\theta,\gamma)=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}(y_{1}-f_{1}y_{2}),$ (11)
$X^{+}(\theta,\gamma)=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}\underline{X},$ (12)
$u^{+}=\Omega(\gamma)^{-1/2}(u_{1}-f_{1}u_{2})$ (13)
and $\operatorname{Var}\left(u^{+}|\mathcal{Z}_{n}^{*}\right)=I_{n}$. In light of (10)-(13) we define
$u^{+}(\theta,\gamma)$ $=y^{+}(\theta,\gamma)-X^{+}(\theta,\gamma)\delta$
$=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}(y_{1}-f_{1}y_{2}-\text{$\underline{X}$$\delta$})$ (14)
and observe that $u^{+}(\theta,\gamma)=u^{+}$ at the true parameters.
The transformation leading to $u^{+}$ consists of three components. Quasi-differencing eliminates classroom and individual effects. The operator $(I+\rho M)^{-1}$ removes cross-sectional correlation and can be understood as a form of spatial Cochrane-Orcutt transformation. Finally, the operator $\Omega(\gamma)^{-1/2}$ scales the spatially uncorrelated residuals to unit variance in the case of heteroskedasticity.
We next provide an outline of our GMM estimation methodology. For clarity we explicitly denote in the following the true parameter vectors as $\theta_{0}=(\rho_{0},f_{1,0},\delta_{0}^{\prime})^{\prime}$ and $\gamma_{0}=(\gamma_{1,0},...,\gamma_{J,0})^{\prime}$. Rigorous consistency and asymptotic normality results as well as additional assumptions needed to establish these properties will be given in Section 4 below. Our GMM estimator is based on both linear and quadratic moment conditions. We consider instrument matrices $H$ and $A$ and require that all elements of $H$ and $A$ are functions of the observable variables that are measurable with respect to $\mathcal{Z}_{n}$. Note that by construction, these variables may depend on the unobserved effects $\alpha$ and $\kappa$. In particular, we set $H=[\underline{X},z]$ as an $n\times q$ instrument matrix.
$m_{n}(\theta,\gamma)=n^{-1/2}\left(\begin{array}[]{c}H^{\prime}u^{+}(\theta,\gamma)\\ u^{+\prime}(\theta,\gamma)Au^{+}(\theta,\gamma)\end{array}\right).$
Under the maintained assumptions of this paper the results in Kuersteiner and Prucha (2020), Theorem 1, imply that the linear moment function is uncorrelated with the quadratic moments. We have
[…]
$Q_{n}\left(\theta,\gamma\right)$ $=\frac{1}{n}m_{n}(\theta,\gamma)^{\prime}\hat{\Xi}_{n}^{-1}m_{n}(\theta,\gamma)$
$=\frac{1}{n}u^{+}(\theta,\gamma)^{\prime}H\left(H^{\prime}H\right)^{-1}H^{\prime}u^{+}(\theta,\gamma)+\frac{1}{n}\frac{\left(u^{+}(\theta,\gamma)^{\prime}Au^{+}(\theta,\gamma)\right)^{2}}{2\operatorname{tr}\left(A^{2}\right)}$ (18)
such that the GMM estimator for $\theta_{0}$ is defined as $\hat{\theta}\left(\gamma\right)=\operatorname{argmin}_{\theta}Q_{n}\left(\theta,\gamma\right)$ for a fixed value of $\gamma.$ In applications $\gamma$ is replaced with a first step consistent estimator. We discuss consistent estimation of $\gamma$ and feasible versions of $\hat{\theta}\left(\gamma\right)$ in Section 4.
@@
A simple two-step procedure to obtain initial estimates for $\delta$, $f_{1}$, and $\rho$ may sometimes be sufficient and can be used to estimate $\gamma.$ Such a two-step procedure can be readily defined adapting ideas from the literature on spatial Cliff-Ord models; cp. e.g., Kelejian and Prucha (1998, 1999). The first step relies on linear moment conditions for the identification of the parameters $\delta$ and $f_{1}$ which exploit that $H$ and the disturbance vector $(I+\rho M)(u_{1}-f_{1}u_{2})$ of model (8) are orthogonal. The idea is to concentrate out $\delta$ and $f_{1}$ in the first stage using the linear moment function. Note that by Lemma 2.1 the linear moment condition $E\left[H^{\prime}\left(y_{1}-y_{2}f_{1}-\underline{X}\delta\right)\right]=0$ holds at the true parameter values for $\delta$ and $f_{1}.$ For simplicity consider the case where $z$ is of dimension $n\times 1$. Then this equation corresponds to a regular just identified linear instrumental variables problem with included exogenous covariates $\underline{X}$ and where $z$ is instrumenting for the endogenous variable $y_{2}$. Letting $W=\left(y_{2},\text{$\underline{X}$}\right)$ and solving the sample analog of the moment condition yields
$\left(\begin{array}[]{c}\tilde{f}_{1}\\ \tilde{\delta}\end{array}\right)=\left(H^{\prime}W\right)^{-1}H^{\prime}y_{1}.$ (19)
@@
In the second step we estimate $\rho$ using a form of the quadratic moment function, where we plug into the residual vector $\tilde{\delta}$ and $\tilde{f}_{1}$ for $\delta$ and $f_{1}$. Let
$\epsilon^{+}\left(\theta\right)=\epsilon^{+}\left(\rho,f_{1},\delta\right)=\left(I+\rho M\right)^{-1}\left(y_{1}-f_{1}y_{2}-\underline{X}\delta\right)$
and consider the quadratic moment vector
$m_{\epsilon}^{q}\left(\rho\right)=\frac{1}{n}\epsilon^{+}\left(\rho,\tilde{f}_{1},\tilde{\delta}\right)^{\prime}A\epsilon^{+}\left(\rho,\tilde{f}_{1},\tilde{\delta}\right)$
with $A$ defined as before. The first stage estimator for $\rho$ then is obtained as

### v2

@@ 3.1 Model with Covariates
Define the sigma fields $\mathcal{F}_{n,i,t}=\sigma\left(S,\alpha,\varphi,X,z,\zeta^{\ast},y_{-t}^{\ast},y_{-i,t}^{*}\right)\text{\,for $t=1,2,$}$ where $z$ denotes variables excluded from $X$ that may be used as instruments. Let $\mathcal{Z}_{n}^{*}=\sigma\left(S,\alpha,\varphi,X,z,\zeta^{\ast}\right)$ be the sigma field of all conditioning variables and let $\mathcal{Z}_{n}=\sigma\left(X,z\right)\subset\mathcal{Z}_{n}^{*}$ be the subset of observable information.
We maintain Assumption 2 without change, but using the expanded versions of $\mathcal{F}_{n,i,t}$, $\mathcal{Z}_{n}^{*}$ and $\mathcal{Z}_{n}$ in Definition 3.1. Then, under Assumptions 2 and 3 the conclusions of Theorem 2.1 continue to hold. Furthermore, we have $\operatorname{Cov}\left(u_{it},X\right)=0$ and $\operatorname{Cov}(u_{it},z)=0$. We note that Assumptions 2, 3 and Theorem 2.1 imply the peer effects model with covariates for test score outcomes
$y_{ct}$ $=\mu_{c}^{*}$ $+v_{c}^{c}\beta_{vt}^{c}+w_{ct}^{c}\beta_{wt}^{c}+\left(I_{c}+\rho M_{c}\right)(v_{c}^{p}\beta_{vt}^{p}+w_{ct}^{p}\beta_{wt}^{p}+u_{ct}).$ (8)
We note that (8) imposes the restriction that average residualized test scores for both tests $t=1$ and $t=2$ are the same. This is in line with our core identification strategy of the two tests testing essentially identical skills.
3.2 Empirical Specification
[…]
$y_{ct}=\mu_{c}^{*}f_{t}+v_{c}^{c}\beta_{vt}^{c}+w_{ct}^{c}\beta_{wt}^{c}+\left(I_{c}+\rho M_{c}\right)(v_{c}^{p}\beta_{vt}^{p}+w_{ct}^{p}\beta_{wt}^{p}+u_{ct}),$ (9)
and where our assumptions imply that $f_{t}=1$ for $t=1,2.$ Empirically we treat $f_{1}$ as a free parameter to be estimated, while we maintain $f_{2}=1$ as a normalization. An estimated value of $f_{1}$ that differs from $f_{1}=1$ is an indication, although not a rigorous specification test, that Assumption 2 is violated in the data.
By quasi-differencing (9), which eliminates $\mu_{c}^{\ast}$, we obtain
@@ 3.2 Empirical Specification
Collecting terms with $f_{1}$ leads to an interpretation of (3.2) where $y_{c2}-v_{c}^{c}\beta_{v2}^{c}-w_{c2}^{c}\beta_{w2}^{c}-\left(I_{c}+\rho M_{c}\right)(v_{c}^{p}\beta_{v2}^{p}+w_{c2}^{p}\beta_{w2}^{p})$ is used to control for unobserved $\mu_{c}^{\ast}$ similar to a control function approach. The difference to a conventional control function approach is that (9) for $t=2$ cannot be consistently estimated by least squares because unobserved components may be correlated with observed covariates, a problem that our GMM estimators address.
We next write (3.2) more compactly by stacking observations across classrooms. Let $v^{p},v^{c}$ and $y_{t},u_{t},w_{t}^{p},w_{t}^{c}$, $t=1,2$, as well as $M$ be as defined above. Furthermore, let $\underline{X}=\left(v^{c},w_{1}^{c},w_{2}^{c},v^{p},w_{1}^{p},w_{2}^{p},Mv^{p},Mw_{1}^{p},Mw_{2}^{p}\right)$ and define
@@
where $\delta_{v}^{c}=\beta_{v1}^{c}-f_{1}\beta_{v2}^{c}$, $\delta_{v}^{p}=\beta_{v1}^{p}-f_{1}\beta_{v2}^{p}$ and $\beta$ is the vector of all $\beta$s above. Then we can write (3.2) for the entire sample more compactly as
$y_{1}=f_{1}y_{2}+\underline{X}\delta+(I+\rho M)(u_{1}-f_{1}u_{2}).$ (12)
@@
where $\delta=\beta_{1}-f_{1}\beta_{2}$. Here $\beta_{t}$ collects the corresponding coefficients for each variable in $\underline{X}$ with elements of $\beta_{t}$ set to 0 if a variable is absent in (9) for given $t$. Let $\mu^{\ast}=(\mu_{1}^{\ast\prime},....,\mu_{C}^{\ast\prime})^{\prime}$. Then we can write (9) compactly as
$y_{t}=\mu^{\ast}f_{t}+\text{$\underline{X}$$\beta_{t}+(I+\rho M)u_{t}$}.$ (13)
@@
With Assumptions 2 and 3, and full column rank conditions on $\underline{X}$, we can identify $\delta$, which is sufficient for the identification of $(\rho,f_{1})$. The parameters $\beta_{w1}^{c}$ and $\beta_{w1}^{p}$ can be directly recovered from $\delta$, and $\beta_{w2}^{c}$ and $\beta_{w2}^{p}$ can be identified when $f_{1}\neq 0$. As is common in pure fixed effect panel settings, the effects of test invariant covariates $v$, in our case $\beta_{v1}^{c},\beta_{v2}^{c}$ and $\beta_{v1}^{p},\beta_{v2}^{p}$ are not identified when using within or differencing estimators. In our case, except when $f_{1}=1,$ we are quasi-differencing the equation and allowing for heterogeneity in parameters between $t=1$ and $t=2.$ This leads to a formulation of the model where $v^{c}$ and $v^{p}$ enter the quasi, or fully differenced equation. However, without further assumptions the parameters $\beta_{v1}^{c}$, $\beta_{v2}^{c}$, $\beta_{v1}^{p}$ and $\beta_{v2}^{p}$ are not separately identified. An example of restrictions where the parameters $\beta_{v1}^{c}$ and $\beta_{v2}^{c}$ are identified arises when $f_{1}\neq 1$ and $\beta_{v1}^{c}=\beta_{v2}^{c}$.
3.3 Estimation
@@
In empirical applications it is sometimes sufficient to estimate $\beta$ or even just $\delta,$ for example when $\underline{X}$ contains contextual peer effects. Conventional regression methods applied directly to (9) are invalid because unobserved fixed effects collected in $\mu_{c}^{\ast}$ may be correlated with $\underline{X}.$ We show in Section 5 that instrumenting for $y_{2}$ in (12) identifies the parameters $\delta$ and $f_{1}.$ We show in Corollary 5.1 that when test scores are recorded as non-negative values the classroom fixed effect or an overall constant is a valid instrument. More generally, exogenous covariates that vary at the classroom level but are excluded from $\underline{X}$ are valid instruments by Lemma 5.1. We consider instrument matrices $H$ and $A$ and require that all elements of $H$ and $A$ are functions of the observable variables that are measurable with respect to $\mathcal{Z}_{n}$. Note that by construction, these variables may depend on the unobserved effects $\alpha$ and $\kappa$. In particular, we set $H=[\underline{X},z]$ as an $n\times q$ instrument matrix.
@@
A simple two-step procedure is sufficient to obtain consistent but inefficient estimates for $\delta$, $f_{1}$, and $\rho$. Such a two-step procedure can be readily defined adapting ideas from the literature on spatial Cliff-Ord models; cp. e.g., Kelejian and Prucha (1998, 1999). The first step relies on linear moment conditions for the identification of the parameters $\delta$ and $f_{1}$ which exploit that $H$ and the disturbance vector $(I+\rho M)(u_{1}-f_{1}u_{2})$ of model (12) are orthogonal. Note that by Theorem 2.1 the linear moment condition $E\left[H^{\prime}\left(y_{1}-y_{2}f_{1}-\underline{X}\delta\right)\right]=0$ holds at the true parameter values for $\delta$ and $f_{1}.$ For simplicity consider the case where $z$ is of dimension $n\times 1$. Then this equation corresponds to a regular just identified linear instrumental variables problem with included exogenous covariates $\underline{X}$ and where $z$ is instrumenting for the endogenous variable $y_{2}$. Letting $W=\left(y_{2},\text{$\underline{X}$}\right)$ and solving the sample analog of the moment condition yields $\left(\tilde{f}_{1},\tilde{\delta}^{\prime}\right)^{\prime}=\left(H^{\prime}W\right)^{-1}H^{\prime}y_{1}$.
@@
In the second step we estimate $\rho$ using a form of the quadratic moment function with, for example, $A=M$. Quadratic moments are motivated by Theorem 2.1 which indicates that the variance-covariance matrix of residual vector $\epsilon^{+}=u_{1}-f_{1}u_{2}$ is diagonal. Hence $E(\epsilon^{+\prime}A\epsilon^{+})=0$ if $A$ has all diagonal elements 0. This suggests plugging $\tilde{\delta}$ and $\tilde{f}_{1}$ into $\epsilon^{+}\left(\rho,f_{1},\delta\right)=\left(I+\rho M\right)^{-1}\left(y_{1}-f_{1}y_{2}-\underline{X}\delta\right)$ and forming the quadratic moment vector $m_{\epsilon}^{q}\left(\rho\right)=\frac{1}{n}\epsilon^{+}\left(\rho,\tilde{f}_{1},\tilde{\delta}\right)^{\prime}A\epsilon^{+}\left(\rho,\tilde{f}_{1},\tilde{\delta}\right)$. The first stage estimator for $\rho$ then is obtained as